Shirley, a recent college​ graduate, excitedly described to her older sister the $1,510 sofa, table, and chairs she found today.​ However, when asked she could not tell her sister which interest calculation method was to be used on her​ credit-based purchase. Calculate the monthly payments and total cost for a bank loan assuming a​one-year repayment period and 13.25 percent interest.​ Now, assume the store uses the​ add-on method of interest calculation. Calculate the monthly payment and total cost with a​ one-year repayment period and 11.25 percent interest. Using the information​ above, how much interest will Shirley​ "save" or be rebated if she can repay the loans after six​ months?

For a bank loan assuming a​ one-year repayment period and 13.25​% ​interest, the monthly payment is?
Answer #1
One year repayment and 13.25% interest
Pv Amount of Purchase $1,510
Rate Monthly interest=(13.25/12)%= 1.10417%
Nper Number of months of payment 12
PMT Monthly Payment $135.05 (using PMT function of excelwith Rate=1.10417%,Nper=12, Pv=-1510)
Excel command:PMT(1.10417%,12,-1510)
Total Cost=135.05*12= $1,620.56
One year repayment and 11.25% interest(ADD ON METHOD)
Amount of Purchase $1,510
Total interest in one year =1510*11.25% $169.88
Total Principal to be paid $1,510
Total Payments=1510+169.88= $1,680
Number of months of payment 12
Monthly Payment=1680/12 $139.99
Total Cost $1,680
Repayment after 6 months at 13.25% interest $1,606.93 1510*(1.1325^0.5)
Saving in interest =1620.56-1606.93 $13.63